OpenStudy (anonymous):

A large asteroid crashed into a moon of a planet, causing several boulders from the moon to be propelled into space toward the planet. Astronomers were able to measure the speed of one of the projectiles. The distance (in feet) that the projectile traveled each second, starting with the first second, was given by the arithmetic sequence 29, 40, 51, 62, . . . Find the total distance that the projectile traveled in seven seconds. 339 feet 405 feet 434 feet 182 feet

OpenStudy (anonymous):

i think its a but i'm not sure...

jimthompson5910 (jim_thompson5910):

use the formula Sn = n(2a1 + d(n-1))/2 where in this case a1 = 29 (first term) d = 11 (common difference) n = 7 (number of terms you want to sum)

OpenStudy (anonymous):

I TRIED TO DERIVE ALL OF MATHEMATICS BUT FAILED...

OpenStudy (anonymous):

I am still very confused...

jimthompson5910 (jim_thompson5910):

Sn = n(2a1 + d(n-1))/2 S7 = 7*(2*29 + 11(7-1))/2 S7 = ???

OpenStudy (anonymous):

434

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Thank you :)

jimthompson5910 (jim_thompson5910):

you're welcome